OpenStudy (anonymous):

Let f (x)=x^3+px^2+qx. Find the values of p and q so that f (-1)=8 and f'(-1)=12

OpenStudy (anonymous):

replace \(x\) by \(-1\) get an equation in \(p\) and\(q\)

OpenStudy (anonymous):

take the derivative, replace \(x\) by \(-1\) get another equation in \(p\) and \(q\) then solve the system of equations

OpenStudy (dayakar):

f (x)=x^3+px^2+qx. Find the values of p and q so that f (-1)=8 and f'(-1)=12 f(-1)=8 \[\left( -1 \right)^{3}+p \times \left( -1 \right) ^{2}+q \times \left( -1 \right)=8\]

OpenStudy (anonymous):

the first equation is \[-1+p-q=8\] or \[p-q=9\] i will let you find the second one

OpenStudy (dayakar):

-1+p-q=8

OpenStudy (anonymous):

3-2p+q=12?

OpenStudy (dayakar):

f(-1) =8 f(-1)= 12 the problem is wrong

OpenStudy (dayakar):

both times you typed f(-1), verify

OpenStudy (anonymous):

Second one was sipped to be f'(-1)=12

OpenStudy (anonymous):

Dealing with derivatives on the first degree